Rod is 15 times as old as Sloan. In 5 years, Rod will be only 5 times as old as Sloan. How old are they now?
Answer by Maths68(1474)

Let
r = current age of Rod
s = current age of Sloan
Rod is 15 times as old as Sloan means
r=15s..........(1)
In 5 years means we will add 5 years in the ages of Rod and Sloan
r+5=s+5
Then Rod will be only 5 times as old as Sloan
r+5=5(s+5)
r+5=5s+25
r=5s+25-5
r=5s+20...........(2)
Put the value of r from (1) to (2)
r=5s+20
15s=5s+20
15s-5s=20
10s=20
s=20/10
s=2
Put the value of s in (1)
r=15s..........(1)
r=15(2)
r=30
Rod is 30 and Sloan is 2 years old